What will be the temp. diff. needed in hot air balloon to lift 1 Kg mass? Assume that volume is 100 m^3, the temp. of air is 25 C, the pressure is 1 bar and air is ideal gas with avg. Molar Mass of 29 g/ mole.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago


hi Vishrant,

let d1 and d2 be the densities of gas inside and outside the ballon

(d2-d1) V g = mg=1 kg *g

d2-d1= 1/100=0.01

d1= d2-0.01 =PM/RT - 0.01 = (10^5 * 29* 10^-3/8.314*298)-0.01

d1=1.17-0.01

   =1.16 kg/m3

T=PM/Rd = 300.1 K